OpenStudy (anonymous):

which of the following expressions represents the product of 3 less than twice x and 2 more the quantity 3 times? a. -6x^2+ 25x + 6 b. 6x^2 +5x + 6 c. 6x^2 - 5x +6 d. 6x^2 - 5x -6 e. 6x^2 - 13x -6

OpenStudy (anonymous):

(2x-3)(3x+2) 6x^2+-9x+4x-6 =6x^2-5x-6

OpenStudy (anonymous):

can you explain it to me please??

OpenStudy (anonymous):

3 less than twice x means 2*x-3

OpenStudy (anonymous):

quatity 3 times, =3x 2 added to it. that is, 3x+2

OpenStudy (anonymous):

(2x-3)*(3x+2)

OpenStudy (anonymous):

just multiply .

OpenStudy (anonymous):

2x(3x+2)-3(3x+2) =6x^2+4x-9x-6 =6x^2-5x-6

OpenStudy (anonymous):

thank you